• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Olson

    Log entry time 08:02:47 on June 3,2007

    Entry number 204131

    keyword=Owl Shift Summary

    Owl Shift Summary:

    M. Olson, E. Hrosinkova

    Shift begins with run 4004 in progress.

    0023: End run 4004 (He4, Q2=63, 53uA, 1.25M)
    0048: Current to 60uA, MCC trying to tune x position on BPM A
    Stopped run 4005 while this is going on.

    0101: Beam repostioned on both BPM x and y. Start run 4006
    0128: Beam current back to 53uA as Hall C has come back on.
    0152: end run 4006 (He4, Q2=63, 60/53uA, 1.25M)
    0157: end run 4007 (90k) to adjust prescales
    0158: start run 4008

    0240: end run 4008 (He4, Q2=63, 53/60uA, ~900K) to adjust prescales
    start run 4009
    0251: end run 4009 to (again) adjust prescales. We are going to leave the prescales at the values for 53uA and not attempt to change them to follw the changes in current. (Too much overhead)
    Start run 4010.

    0354 end run 4010 (He4, Q2=63, 55uA, 1.25M)
    start run 4011

    0438: end run 4011 (He4, Q2=63, 55/45uA, 705k)
    adjust prescales for 45uA
    start run 4012

    0550 end run 4012 (He4, Q2=63, 45-60uA, 1.25M)
    Many beam trips and 10 min beam down during run

    0657: end run 4013 (He4, Q2=63, 60/50uA, 1.25M)
    start run 4014

    Shift end with 4014 in progress